OpenStudy (anonymous):

HELP! Solve the following system of equations. Enter the y-coordinate of the solution. Round your answer to the nearest tenth. 5x+2y=21 -2x+6y=-34

OpenStudy (tkhunny):

How do you propose to solve these? Personally, I would multiply the first by 2 and the second by 5 and then add them.

OpenStudy (anonymous):

x=5.70 y=-3.76

OpenStudy (anonymous):

I dont know how to go about solving this. This is like foreign to me @tkhunny :(

OpenStudy (tkhunny):

Did you do what I said?

OpenStudy (anonymous):

@giligails so if its asking for just the y coordinate and to round to the nearest tenth would it be 4?

OpenStudy (anonymous):

@tkhunny which numbers should i be multiplying?

OpenStudy (kenljw):

Multiply the equations so either the coefficient of x or y are the same, the eliminate one variable and solve. Place value in initial equation to fine other variable.

OpenStudy (tkhunny):

5x+2y=21 -2x+6y=-34 Multiply the first by 2 10x+4y=42 Multiply the second by 5 -10x+30y=-170 Now add the two equations. 10x - 10x = ?? 4y + 30y = ?? 42 - 170 = ??

OpenStudy (anonymous):

first 2y=21-5x therefore -2x+3(21-5x)=-34 -2x-15x=-34-63 -17x=-97 x=97/17 2y=21-97*5/17 21-28.529=2y 2y=-7.529 y=-3.764

OpenStudy (tkhunny):

Now that I think about it, I would have left the first alone and multiplied the second by 5/2. No matter. It will all work out if no error is made.

OpenStudy (anonymous):

@tkhunny thank you, so how would those steps be broken down? if you dont mind

OpenStudy (anonymous):

@giligails so the answer is simply -3.764? is that including rounding it up to the nearest tenth?

OpenStudy (tkhunny):

Not a good question. Those ARE the steps broken down. Did you do the additions? There are three of them there for you.

OpenStudy (anonymous):

I got 0, 34y and -128. but i dont know how which one would be the answer @tkhunny

OpenStudy (tkhunny):

You are adding equations. You should be left with an equation. 0 + 34y = -128 or just 34y = -128 Solve for y.